(x+3)^3 this should be clubed. factor and give result please. i beleve that the book is wrong. it says the answer is x squared plus 6x plus 9.
Answer by stanbon(48568)

(x+3)^3 = x^3 + 3x^2(3) + 3x(3)^2 + 3^3
=x^3 + 9x^2 27x +27
------
Cheers,
Stan H.